Good day! Can you help me solve this question?
The areas of two similar tetrahedrons are 24 sq. in and 96 sq. in., find the ratio of their volumes.

Thank you!

Answer by CubeyThePenguin(3113)

area ratio = 24/96 = 1/4

side ratio = sqrt(1/4) = 1/2

volume ratio = (1/2)^3 = 1/8